OpenStudy (anonymous):

What is the correct simplified form for this porblem with steps so i can understand fully? (2)/(x+3)-(1)/(x-3)+(2x)/(x²-9)

OpenStudy (compassionate):

Is this for school? If not I can just show you all the steps.

OpenStudy (harsimran_hs4):

\[\frac{ 2 }{ x+3 } - \frac{ 1 }{ x+3 } + \frac{ 2 }{ (x+3) (x-3)}\] can you take common denominator now?

OpenStudy (anonymous):

its for school but we are allowed to use this my whole class was spose to make one

OpenStudy (anonymous):